rssLink RSS for all categories
 
icon_red
icon_green
icon_red
icon_red
icon_red
icon_green
icon_green
icon_red
icon_red
icon_red
icon_orange
icon_green
icon_green
icon_green
icon_red
icon_blue
icon_red
icon_orange
icon_red
icon_red
icon_red
icon_red
icon_red
icon_red
icon_red
icon_green
icon_red
icon_green
icon_green
 

FS#5678 — FS#9596 — var-1-6k

Attached to Project— Network
Incident
the whole network
CLOSED
100%
The BGP session with UPC is unstable at certain times of the day. We have effectively detected an anomaly on the router at the time of the automatic daily backup of the configuration.
Date:  Tuesday, 21 January 2014, 10:41AM
Reason for closing:  Done
Comment by OVH - Monday, 28 October 2013, 13:16PM

The session with UPC is currently unstable and access to the configuration is no longe possible due to a problem on the NVRAM. We will reboot the router as a matter of urgency.


Comment by OVH - Monday, 28 October 2013, 14:22PM

The router reboot is in progress. Tarffic towards Netia, UPC PL and TPIX has been rerouted towards Frankfurt and PLIX.


Comment by OVH - Monday, 28 October 2013, 14:23PM

The supervisor has crashed. We are trying to reboot again.


Comment by OVH - Monday, 28 October 2013, 15:32PM

The router is stable. We are restoring the traffic.


Comment by OVH - Monday, 28 October 2013, 15:33PM

The router is on line but the BGP sessions remain unstable. We will replace the supervisor.


Comment by OVH - Monday, 28 October 2013, 17:21PM

We have identified a section of the config that hadn't transferred correctly onto the router following the chassis reboots. We have fixed it but now we have another problem with the OSPF. We're investigating.


Comment by OVH - Tuesday, 29 October 2013, 01:22AM

The router is now up and the OSPF is mounted.

-We are having instabilities on the BGP session with UPC: we are keeping them in touch.
-The BGP session with the TPIX route servers are down: we are keeping them in touch too.


Comment by OVH - Tuesday, 29 October 2013, 01:23AM

We have redirected the traffic to UPC PL in Frankfurt to avoid the saturation of our peering in Amsterdam.


Comment by OVH - Thursday, 31 October 2013, 07:22AM

The router has rebooted but it seems that there's an electrical issue this time. We are again having the OSPF issue.


Comment by OVH - Thursday, 31 October 2013, 17:04PM

We're going to change the router supervisor.


Comment by OVH - Thursday, 31 October 2013, 17:04PM

We're shutting down the router.


Comment by OVH - Thursday, 31 October 2013, 17:05PM

The router rebooted. We have to update it.


Comment by OVH - Thursday, 31 October 2013, 17:06PM

The router is up. We are reconfiguring it, then we will restore the traffic.


Comment by OVH - Thursday, 31 October 2013, 17:19PM

There is some instability on the OSPF. We're investigating.


Comment by OVH - Monday, 04 November 2013, 14:49PM

The router has been stable since Friday but there are still some anomalies and we continue to work on the case with Cisco. Peerings towards TPIX are still deactivated for the time being. The anomalies are on OSPF protocol (OSPF adjacencies take a very long time to set up following a reboot of the chassis) and BGP level (some sessions are unstable, including those with the TPIX Route-Server, some peers have brought up abnormal route updates).